OpenStudy (torialaw192):

An inequality is shown below: −np − 4 ≤ 2(c − 3) Which of the following solves for n? n ≤ − the quantity 2 times c minus 2 all over p n ≤ − the quantity 2 times c minus 10 all over p n ≥ − the quantity 2 times c minus 2 all over p n ≥ − the quantity 2 times c minus 10 all over p

sam (.sam.):

You've got to isolate 'n', do you know how to do that?

OpenStudy (torialaw192):

@.Sam. Nu..

OpenStudy (torialaw192):

@pooja195 Can you help me please?

OpenStudy (welshfella):

first isolate -np by adding 4 to both sides of the inequality -np -4 + 4 <= 2(c - 3) + 4 can you simplify that?

OpenStudy (welshfella):

-4 + 4 = ?

OpenStudy (torialaw192):

0

OpenStudy (welshfella):

ok so we have -np + 0 <= 2(c - 3) + 4 the 0 disappears . so -np <= 2(c - 3) + 4 now we divide both sides by -p to get n NOte when you divide by a a negative with these inequalities the sign is flipped and so here it becomes >= .

OpenStudy (welshfella):

p >= [2(c - 3) + 4] / -p if you simplify that you'll get the correct choice

OpenStudy (welshfella):

* that should be n >=

OpenStudy (torialaw192):

Ok.